2*3 consecutive even numbers such as the sum of twice the first number and the third number is 30 more than the second number.
equation__________________
Answer____________________
how do you do this? thank you for your help. i have 5 more to do after this and i am confused.

Answer by edjones(7311)

let the 3 integers be: x, x+2, x+4
2x+x+4=30+x+2
3x+4=x+32
subtract x and 4 from each side: 2x=28
x=14, x+2=16, x+4=18
Check:
2*14+18=30+16
28+18=46
Hope that helps.
Ed